OpenStudy (lovelyanna):

What algebraic property does this statement show? 3 + (–7) = (–7) + 3 A. associative property B. symmetric property C. commutative property D. closure property Is it C?

OpenStudy (anonymous):

Yes it is. :-)

OpenStudy (lovelyanna):

Thanks :)

OpenStudy (lovelyanna):

Can you help with one more?

OpenStudy (lovelyanna):

@ChristopherToni

OpenStudy (anonymous):

Sure. Ask away! :-)

OpenStudy (lovelyanna):

Or D? :/

OpenStudy (anonymous):

Unfortunately it's not (A). What you want to do is first draw the line 2x+3y = 6. In slope intercept form this would be \(\large y=-\frac{2}{3}x + 6\). So the line you're working with is correct (the < part implies the line is dashed, not solid; hence the answer can't be (B) or (C)). Now, what you want to do is pick a point above and below the line and plug it into the inequality \(2x+3y < 6\). If the point above the line satisfies the inequality, you shade above the line, and if the point below the line satisfies the inequality, you shade below the line. With that said, it's not (A) because if you take the point (0,3) which is in the region above the line, you get \(2x+3y = 2(0)+3(3) = 9 \not< 6\), where as if you take the point (0,0) which is below the line, you get \(2x+3y = 2(0)+3(0) = 0 < 6\). Therefore, the region you want to shade is below the line, and hence the correct answer is (D). I hope this makes sense! :-)

OpenStudy (lovelyanna):

Thought so, then this would be D, that was my second answer :P Very long answer, and very informative. Thank you very much!!

OpenStudy (lovelyanna):

Emily, you here?

jigglypuff314 (jigglypuff314):

hi :)

OpenStudy (lovelyanna):

I was waiting for you to see if #1: C and #2: D are correct, are they?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Oh, ok, thanks :)

OpenStudy (lovelyanna):

I gtg now, thanks for your help! ♥